OpenStudy (anonymous):

solve the equasion : sinx/cosx = 0

OpenStudy (anonymous):

when sin(x) = -cos(x) which is in the second and fourth quadrants at a 45deg angle or deg 135 or 315.

OpenStudy (raden):

sinx/cosx can rewriten as tanx so, tanx = 0 what angles make the tan be 0 ? these is special angles in trigonometry

OpenStudy (anonymous):

thank you

OpenStudy (raden):

you're welcome, i sure u got it :)

OpenStudy (anonymous):

yes i understood it.but hey,what about the equasion : sinx/x = 0

OpenStudy (raden):

sinx/x ?

OpenStudy (anonymous):

well yes.sinx devided by x = 0 and we have to solve the equasion

OpenStudy (raden):

i think we can cross multiplied, because it is equation = 0 so, sinx/x = 0 sinx = x * 0 sinx = 0 solve for x

OpenStudy (anonymous):

uhm,okk,thanks again

OpenStudy (raden):

welcome again :D